Aaron Purr, the best cat in the world, chases an average of 12 birds per day. He is not a mighty hunter and rarely catches any. But that's ok, we'll use his chasing info for this problem. a. What is the probability that, today, Aaron Purr will chase more than 9 birds? b. What is the probability that Aaron Purr will chase exactly 7 birds within 12 hours tomorrow?

PLEASE SHOW DETAILED WORK ON EACH ANSWER
Answer #1

Answer:-

Given That:-

Aaron Purr, the best cat in the world, chases an average of 12 birds per day. He is not a mighty hunter and rarely catches any. But that's ok, we'll use his chasing info for this problem.

Given,

X : Number of birds catched by Aaron Purr per day average x = 12

Here chance of catching a bird is very high but accurance is less.

Catched birds are independent of each other

Hence

X\sim Poisson(\lambda=12)

a. What is the probability that, today, Aaron Purr will chase more than 9 birds?

P(X > 9) = 1-P(X\leq 9)

=1-\sum_{x=0}^{9}\farc{e^{-12}(12)^x}{x!}

= 1 - 0.2424

= 0.7576

b. What is the probability that Aaron Purr will chase exactly 7 birds within 12 hours tomorrow?

Y ~ Number of birds catched in 12 years

Y\sim poisson(\lambda_1=6)

P(Y = 7) =  \frac{e^{-6}(6)^7}{7!}

= 0.1377

Thank you for your supporting. Please upvote my answer...